OpenStudy (anonymous):

Mercury has a density of 5.300 g/mL. How much would 10.0 g of mercury occupy?

OpenStudy (anonymous):

d=m/V 5.3g/mL=10g/xmL xmL=10g/(5.3g/mL) xmL=1.89 mL?

OpenStudy (dumbcow):

density = mass/volume plug in what you know

OpenStudy (anonymous):

OH SORRY. it's not 5.300, it's 5.00

OpenStudy (anonymous):

never mind i go it. thank you guys! (: